Temple University Philadelphia, Pennsylvania 19122-2595 Department of Mathematics e-mail: zeilberg@math.temple.edu

(215) 204–7287 (voice) (215) 204–6433 (Fax)

Jan. 22, 2001

Professor Stephen J. Greenfield Department of Mathematics Rutgers University New Brunswick, NJ 08901

Dear Steve,

This letter is in support of Mohamud Mohammed, in support of his application to the graduate program in the mathematics department at Rutgers.

Mohammed is currently my Ph.D. student, and I very much hope that he will be able to transfer with me to Rutgers. He currently has a research assistanship, paid by my current NSF grant, that is scheduled to expire June 2001. In my pending grant I have asked for a continuation. If this materializes, then Mohamud will not need a TAship, but if it does not, it would be nice if you would be also able to offer him support.

I am very pleased with Mohamud, who is one of my best students ever. He has a very solid background in traditional mathematics, and at the same time is becoming a true computer-math whiz in the kind of math that I am doing with computers. He has already made significant progress towards his very challenging thesis problem, that of studying the so-called WZ cohomology.

Mohamud has subbed for me several times in both my Freshman Calculus class and advanced computer-algebra, and has done a great job. So he is also a very good teacher. I very much hope that he will be able to transfer to the Rutgers math program, surely for convenience reason, but also because the stimulating environment at Rutgers will help him grow professionally.

If you have any further questions, I will be glad to answer them. My phone numbers are (609)921-7873 (home), (215)204-7287 (office), and my e-mail address is zeil-

berg@math.temple.edu .

Sincerely,

Doron Zeilberger Professor of Mathematics